# Analyzing Multifiltering Functions Using Multiparameter Discrete Morse Theory

### v1

@@ 1. Introduction
These ideas are pushed a step further to the multiparameter framework by Allili et al. [3]. In their paper, the authors prove that a discrete gradient field, also called a Morse matching, can be used to reduce a Lefschetz complex, e.g. a simplicial or cubical complex, while preserving multipersistent homology. Moreover, they present the algorithm Matching that takes as input a simplicial complex and a vector function defined on its vertices and outputs a Morse matching compatible with the multifiltration induced by the given multifiltering function. By restraining the considered input functions to component-wise injective vertex maps, the algorithm Matching is improved in [4] and a more efficient version, ComputeDiscreteGradient, is proposed thereafter by Scaramuccia et al. [61].
@@
The paper is structured as follows. In Section 2, we recall some known concepts, mainly about multipersistent homology. In Section 3, a brief overview of mdm theory is presented. Then, our main algorithm GenerateMDM and subroutines are described in Section 4. We analyze its complexity and compare it with previous similar algorithms Matching and ComputeDiscreteGradient. Next, we prove in Section 5 that GenerateMDM outputs the desired result. Thereafter, the notion of discrete Pareto set is introduced in Section 6. Finally, in Section 7, we see how GenerateMDM performs in practice and we experiment with the concept of discrete Pareto set and that of critical components of a mdm function on triangular meshes.
@@ 2.1. Simplicial complexes
Moreover, the set $K$ may be seen as a topological space by endowing it with the Alexandrov topology [7, 53, 64]. A set of simplices $S\subset K$ is open in this topology if $\tau\geq\sigma\in S$ implies that $\tau\in S$. Hence, the smallest neighbourhood of a simplex $\sigma\in K$ is its star $\operatorname{St}\sigma=\{\tau\in K\ |\ \tau\geq\sigma\}$. Also, it could be shown that for any $S\subseteq K$, its closure is $\operatorname{Cl}S=\{\tau\in K\ |\ \tau\leq\sigma\text{ for some }\sigma\in S\}$ and its interior is $\operatorname{Int}S=\{\sigma\in S\ |\ \tau\in S\text{ for all }\tau\geq\sigma\}$. Using these sets, we can define the boundary and the exit set of $S$ as $\operatorname{Bd}S:=\operatorname{Cl}S\backslash\operatorname{Int}S$ and $\operatorname{Ex}S:=\operatorname{Cl}(S)\backslash S$ respectively.
@@ 2.2. Homology
Therefore, we have $\beta_{0}(K)=1$, $\beta_{1}(K)=1$ and $\beta_{2}(K)=0$ when considering homology with coefficients in ${\mathbb{Z}}$, but $\beta_{0}(K;{\mathbb{Z}}_{2})=1$, $\beta_{1}(K;{\mathbb{Z}}_{2})=2$ and $\beta_{2}(K;{\mathbb{Z}}_{2})=1$ when using coefficients in ${\mathbb{Z}}_{2}$. The ${\mathbb{Z}}_{2}$ summand in $H_{1}(K)$ corresponds to the torsion of the simplicial complex, which represents in this case the non-orientability of the Klein bottle. This information is lost when considering coefficients in ${\mathbb{Z}}_{2}$. Thus, throughout this paper, we use homology with coefficients in ${\mathbb{Z}}$, unless clearly stated otherwise (as in Table 2).
@@ 2.3. Multiparameter persistent homology
Let $\preceq$ be the coordinate-wise order on ${\mathbb{R}}^{k}$, i.e. for $u,u^{\prime}\in{\mathbb{R}}^{k}$, we have $u\preceq u^{\prime}$ iff $u_{i}\leq u_{i}^{\prime}$ for each $i=1,...,k$. We also note $u\precneqq u^{\prime}$ when $u\preceq u^{\prime}$ and $u\neq u^{\prime}$. A simplicial complex $K$ is multifiltered when it is considered together with a multifiltration, which is a finite family of subcomplexes $\{K(u)\}_{u\in{\mathbb{R}}^{k}}$ of $K$ such that $u\preceq u^{\prime}$ implies $K(u)\subseteq K(u^{\prime})$. An example of a bifiltration, i.e. a multifiltration with $k=2$, is illustrated in Figure 1a.
@@
Moreover, each inclusion $K(u)\hookrightarrow K(u^{\prime})$ induces a linear map $\iota_{p}^{u,u^{\prime}}:H_{p}(K(u))\rightarrow H_{p}(K(u^{\prime}))$ on the associated homology modules for each dimension $p$. Essentially, the image of a map $\iota_{p}^{u,u^{\prime}}$ represents the homology classes of $K(u)$ of dimension $p$ still alive in $K(u^{\prime})$. We define the $p^{\text{th}}$ multiparameter persistence module of a given multifiltration as the family of all homology modules $H_{p}(K(u))$ together with the linear maps $\iota_{p}^{u,u^{\prime}}:H_{p}(K(u))\rightarrow H_{p}(K(u^{\prime}))$. Such a module is depicted in Figure 1b. Multiparameter persistence could be summarized as the study of the multiparameter persistence modules of a multifiltration $\{K(u)\}_{u\in{\mathbb{R}}^{k}}$ or, put simply, of the homological changes in $K(u)$ as the multiple parameters $u_{1},...,u_{k}$ increase.
In this paper, we are mostly interested in multifiltrations induced by a map $f:K\rightarrow{\mathbb{R}}^{k}$. More precisely, we consider the collection of sublevel sets $K(u):=\{\sigma\in K\ |\ f(\sigma)\preceq u\}$. It is easy to see that $K(u)\subseteq K(u^{\prime})$ when $u\preceq u^{\prime}$. However, for $K(u)$ to be a subcomplex of $K$ for all $u$, it is necessary and sufficient to assume that $\sigma<\tau$ implies $f(\sigma)\preceq f(\tau)$, which leads to the following definition.
@@
We call $f:K\rightarrow{\mathbb{R}}^{k}$ an admissible map, or a (multi)filtering function, if $f(\sigma)\preceq f(\tau)$ for all pairs $\sigma<\tau$ in $K$.
@@
In previous related papers [4, 42, 61], the definition of admissible maps was more specific. Namely, a vertex map $f:K_{0}\rightarrow{\mathbb{R}}^{k}$ was considered and assumed to be component-wise injective, i.e. each $f_{i}$ is injective. Then, $f$ was extended to a multifiltering function on all $\sigma\in K$ by setting $f_{i}(\sigma):=\max_{v\in\sigma}f_{i}(v)$ for each $i=1,...,k$. Here, we call $f$ a max-extension of a vertex map if $f_{i}(\sigma)=\max_{v\in\sigma}f_{i}(v)$ for all $i=1,...,k$ and $\sigma\in K$. Although this type of map is very convenient, especially when working with point data, we will see in the examples below that some common filtering functions are not max-extensions. Hence, our broader definition of admissible maps makes it possible to apply the algorithms of this paper to a wider range of input.
Finally, notice that for all $\sigma,\tau\in K$, we have $f(\sigma)\preceq f(\tau)$ iff $f_{i}(\sigma)\leq f_{i}(\tau)$ for each $i=1,...,k$. Therefore, $f$ is admissible iff each $f_{i}:K\rightarrow{\mathbb{R}}$ also is, so we can construct multifiltering function by combining multiple (real-valued) filtering functions. Thus, we end this subsection with a few examples of real-valued maps which may be combined to obtain a multifiltering function:
@@
  • Let $K$ be the Delaunay triangulation of a set of points $K_{0}\subset{\mathbb{R}}^{n}$ and consider its filtration in alpha complexes $\{\mathrm{Alpha}_{r}(K_{0})\}_{r\in{\mathbb{R}}}$. Then, for $\sigma\in K$, note $r_{\sigma}$ the minimal radius parameter $r$ such that $\sigma\in\mathrm{Alpha}_{r}(K_{0})$. The sublevel set $K(r)$ of the radius map $\sigma\mapsto r_{\sigma}$ is exactly $\mathrm{Alpha}_{r}(K_{0})$ for all $r\in{\mathbb{R}}$, meaning that we can see the filtration $\{\mathrm{Alpha}_{r}(K_{0})\}_{r\in{\mathbb{R}}}$ as the sublevel set of the radius map $\sigma\mapsto r_{\sigma}$. A similar reasoning is also valid for point sets filtered in Čech and Vietoris-Rips complexes.
@@ 2.4. Combinatorial vector fields
To ease the computation of multipersistent homology, we can use combinatorial vector fields. Intuitively, we can see a combinatorial vector field on a simplicial complex $K$ as a partition of $K$ into singletons and pairs $(\sigma,\tau)$ such that $\sigma$ is a facet of $\tau$. More formally, a discrete vector field, or a combinatorial vector field, on a simplicial complex $K$ is an injective partial self-map ${\mathcal{V}}:K\nrightarrow K$ such that
@@ 3. $\mathrm{(3)}$
A combinatorial vector field ${\mathcal{V}}$ is said to be acyclic if all its nontrivial ${\mathcal{V}}$-path do not loop, i.e. for any sequence of simplices
$\sigma_{0}^{(p)},\tau_{0}^{(p+1)},\sigma_{1}^{(p)},\tau_{1}^{(p+1)},\sigma_{2}^{(p)},...,\tau_{n-1}^{(p+1)},\sigma_{n}^{(p)}$
in $K$ such that $\sigma_{i}\in\operatorname{dom}{\mathcal{V}}$, ${\mathcal{V}}(\sigma_{i})=\tau_{i}$ and $\tau_{i}>\sigma_{i+1}\neq\sigma_{i}$ for each $i=0,...,n-1$ (where $n\geq 1$), we have $\sigma_{n}\neq\sigma_{0}$.
@@
Consider a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$ on a simplicial complex $K$. We say a combinatorial vector field ${\mathcal{V}}$ is compatible with $f$ or $f$-compatible if $f(\sigma)=f({\mathcal{V}}(\sigma))$ for all $\sigma\in\operatorname{dom}{\mathcal{V}}$.
@@ 3.1. Main definitions
$\mathcal{H}_{g}(\sigma)$ =$=\left\{\tau^{(p+1)}>\sigma\ |\ g(\tau)\preceq g(\sigma)\right\};$
@@ 3. $\mathrm{(3)}$
    For all cofacet $\tau^{(p+1)}>\sigma$, either $g(\tau)\preceq g(\sigma)$ or $g(\tau)\succneqq g(\sigma)$.
@@ 4. $\mathrm{(4)}$
${\mathcal{V}}(\sigma)=\begin{cases}\sigma&\text{ if $\sigma$ is critical},\\ \tau&\text{ if }\mathcal{H}_{g}(\sigma)=\{\tau\}\text{ for some }\tau>\sigma,\\ \text{undefined}&\text{ if }\mathcal{T}_{g}(\sigma)\neq\emptyset.\end{cases}$
@@
Notice that for a mdm function $g$, we have $\mathcal{H}_{g}(\sigma)=\{\tau\}$ iff $\mathcal{T}_{g}(\tau)=\{\sigma\}$. Hence, the gradient field ${\mathcal{V}}$ of $g$ is such that ${\mathcal{V}}(\sigma)=\tau$ iff ${\mathcal{V}}(\tau)$ is undefined and ${\mathcal{V}}^{-1}(\tau)=\sigma$, and the critical simplices of $g$ are exactly the fixed points ${\mathcal{V}}$. Thus, we often refer to the fixed points of a gradient field ${\mathcal{V}}$ as the critical simplices of ${\mathcal{V}}$ instead. Moreover, we could show that a discrete vector field is acyclic if and only if it is the gradient field of some mdm function.
Furthermore, it was shown by Allili et al. [4] that if $f:K\rightarrow{\mathbb{R}}^{k}$ is an admissible map which is the max-extension of a component-wise injective vertex map, there exists a $f$-compatible mdm function. One of the main result of this paper, Theorem 5.7, extends this result. Indeed, we will show that for any admissible map $f:K\rightarrow{\mathbb{R}}^{k}$, we can algorithmically define a mdm function $g:K\rightarrow{\mathbb{R}}^{k}$ which, in addition to being $f$-compatible, is obtained from small perturbations of $f$, and is therefore as close as desired to the initial input.
@@ 3.2. Morse inequalities and (relative) perfectness
When the equality holds for each $u\in f(K)$ and each $p\in{\mathbb{Z}}$, we say ${\mathcal{V}}$ is relative-perfect or perfect relatively to $f$. The concept was first introduced by Fugacci et al. [31] for real-valued filtering functions.
@@
Consider the triangulated circle $K$ embedded in ${\mathbb{R}}^{2}$ as in Figure 3 et let $f:K\rightarrow{\mathbb{R}}^{2}$ be the max-extension of the map which associates each vertex to its coordinates $(x,y)\in{\mathbb{R}}^{2}$. In this particular case, $f$ happens to be a mdm function, and its gradient field is as in Figure 3. The Betti numbers of the circle are $\beta_{0}(K)=\beta_{1}(K)=1$, so $f$ has too many critical simplices to be a perfect mdm function. Nonetheless, it is relative-perfect. Indeed, if $u\in{\mathbb{R}}^{k}$ is such that $m_{p}(u)=0$ for each $p$, meaning that $L_{u}$ does not contain any critical simplex, then we trivially have the equality $m_{p}(u)=\operatorname{rank}H_{p}\left(K(u),\bigcup_{u^{\prime}\precneqq u}K(u^{\prime})\right)=0$. Otherwise, if $u\in{\mathbb{R}}^{k}$ is such that $m_{p}(u)>0$ for some $p$, we can check that $L_{u}$ contains a single critical simplex $\sigma$ of index $p\in\{0,1\}$, i.e. either a critical vertex or a critical edge. We could then compute that $H_{p}\left(K(u),\bigcup_{u^{\prime}\precneqq u}K(u^{\prime})\right)\cong{\mathbb{Z}}$, so $m_{p}(u)=\operatorname{rank}H_{p}\left(K(u),\bigcup_{u^{\prime}\precneqq u}K(u^{\prime})\right)=1$.

### v3

@@ 1. Introduction
These ideas are pushed a step further to the multiparameter framework by Allili et al. [3]. In their paper, the authors prove that a discrete gradient field, also called a Morse matching, can be used to reduce a Lefschetz complex, e.g. a simplicial or cubical complex, while preserving multipersistent homology. Moreover, they present the algorithm Matching that takes as input a simplicial complex and a vector function defined on its vertices and outputs a Morse matching compatible with the multifiltration induced by the given multifiltering function. By restraining the considered input functions to component-wise injective vertex maps, the algorithm Matching is improved in [4] and a more efficient version, ComputeDiscreteGradient, is proposed thereafter by Scaramuccia, Iuricich, De Floriani and Landi [63].
@@
The paper is structured as follows. In Section 2, we recall some known concepts, mainly about multipersistent homology. In Section 3, a brief overview of mdm theory is presented. Then, our main algorithm GenerateMDM and subroutines are described in Section 4. We analyze its complexity and compare it with previous similar algorithms Matching and ComputeDiscreteGradient. Thereafter, the notion of discrete Pareto set is introduced in Section 5. Finally, in Section 6, we see how GenerateMDM performs in practice and we experiment with the concept of discrete Pareto set and that of critical components of a mdm function on triangular meshes.
@@ 2.1. Simplicial complexes
Moreover, the set $K$ may be seen as a topological space by endowing it with the Alexandrov topology [7, 55, 66]. A set of simplices $S\subset K$ is open in this topology if $\tau\geq\sigma\in S$ implies that $\tau\in S$. Hence, the smallest neighbourhood of a simplex $\sigma\in K$ is its star $\operatorname{St}\sigma=\{\tau\in K\ |\ \tau\geq\sigma\}$. Also, it could be shown that for any $S\subseteq K$, its closure is $\operatorname{Cl}S=\{\tau\in K\ |\ \tau\leq\sigma\text{ for some }\sigma\in S\}$ and its interior is $\operatorname{Int}S=\{\sigma\in S\ |\ \tau\in S\text{ for all }\tau\geq\sigma\}$. Using these sets, we can define the boundary and the exit set of $S$ as $\operatorname{Bd}S:=\operatorname{Cl}S\backslash\operatorname{Int}S$ and $\operatorname{Ex}S:=\operatorname{Cl}S\backslash S$ respectively.
@@ 2.2. Homology
Therefore, we have $\beta_{0}(K)=1$, $\beta_{1}(K)=1$ and $\beta_{2}(K)=0$ when considering homology with coefficients in ${\mathbb{Z}}$, whereas $\beta_{0}(K;{\mathbb{Z}}_{2})=1$, $\beta_{1}(K;{\mathbb{Z}}_{2})=2$ and $\beta_{2}(K;{\mathbb{Z}}_{2})=1$ when using coefficients in ${\mathbb{Z}}_{2}$. The ${\mathbb{Z}}_{2}$ summand in $H_{1}(K)$ corresponds to the torsion of the simplicial complex, which represents in this case the non-orientability of the Klein bottle. This information is lost when considering coefficients in ${\mathbb{Z}}_{2}$. Thus, throughout this paper, we use homology with coefficients in ${\mathbb{Z}}$, unless clearly stated otherwise (as in Table 2).
@@ 2.3. Multiparameter persistent homology
Let $\preceq$ be the coordinate-wise order on ${\mathbb{R}}^{k}$, i.e. for $u,u^{\prime}\in{\mathbb{R}}^{k}$, we have $u\preceq u^{\prime}$ iff $u_{i}\leq u_{i}^{\prime}$ for each $i=1,...,k$. We also note $u\precneqq u^{\prime}$ if $u\preceq u^{\prime}$ and $u\neq u^{\prime}$. A simplicial complex $K$ is multifiltered when it is considered together with a multifiltration, which is a finite family of subcomplexes $\{K(u)\}_{u\in{\mathbb{R}}^{k}}$ of $K$ such that $u\preceq u^{\prime}$ implies $K(u)\subseteq K(u^{\prime})$. An example of a bifiltration, i.e. a multifiltration with $k=2$, is illustrated in Figure 1a.
@@
Moreover, for every dimension $p$, each inclusion $K(u)\hookrightarrow K(u^{\prime})$ induces a linear map $\iota_{p}^{u,u^{\prime}}:H_{p}(K(u))\rightarrow H_{p}(K(u^{\prime}))$ on the associated homology modules. Essentially, the image of a map $\iota_{p}^{u,u^{\prime}}$ represents the homology classes of $K(u)$ of dimension $p$ still alive in $K(u^{\prime})$. We define the $p^{\text{th}}$ multiparameter persistence module of a given multifiltration as the family of all homology modules $H_{p}(K(u))$ together with the linear maps $\iota_{p}^{u,u^{\prime}}:H_{p}(K(u))\rightarrow H_{p}(K(u^{\prime}))$. Such a module is depicted in Figure 1b. Multiparameter persistence could be summarized as the study of the multiparameter persistence modules of a multifiltration $\{K(u)\}_{u\in{\mathbb{R}}^{k}}$ or, put simply, of the homological changes in $K(u)$ as the multiple parameters $u_{1},...,u_{k}$ increase.
In this paper, we are mostly interested in multifiltrations induced by a vector map. More precisely, given $f:K\rightarrow{\mathbb{R}}^{k}$, we consider the collection of sublevel sets $K(u):=\{\sigma\in K\ |\ f(\sigma)\preceq u\}$. It is easy to see that $K(u)\subseteq K(u^{\prime})$ when $u\preceq u^{\prime}$. However, for $K(u)$ to be a subcomplex of $K$ for all $u$, it is necessary and sufficient to assume that $\tau<\sigma$ implies $f(\tau)\preceq f(\sigma)$, which leads to the following definition.
@@
We call $f:K\rightarrow{\mathbb{R}}^{k}$ an admissible map, or a (multi)filtering function, if $f(\tau)\preceq f(\sigma)$ for all pairs $\tau<\sigma$ in $K$.
@@
In previous related papers [4, 43, 63], the definition of admissible maps was more specific. Namely, a vertex map $f:K_{0}\rightarrow{\mathbb{R}}^{k}$ was considered and assumed to be component-wise injective, i.e. each $f_{i}$ had to be injective. Then, $f$ was extended to a multifiltering function on all $K$ by setting $f_{i}(\sigma):=\max_{v\in\sigma}f_{i}(v)$ for each $\sigma\in K$ and $i=1,...,k$. Here, we call $f$ a max-extension of a vertex map if $f_{i}(\sigma)=\max_{v\in\sigma}f_{i}(v)$ for all $\sigma\in K$ and $i=1,...,k$. Although this type of map is very convenient, especially when working with point data, we will see in the examples below that some common filtering functions are not max-extensions. Hence, our broader definition of admissible maps makes it possible to apply the algorithms of this paper to a wider range of input.
Finally, notice that for all $\sigma,\tau\in K$, we have $f(\tau)\preceq f(\sigma)$ iff $f_{i}(\tau)\leq f_{i}(\sigma)$ for each $i=1,...,k$. Therefore, $f$ is admissible iff each $f_{i}:K\rightarrow{\mathbb{R}}$ also is, so we can construct multifiltering function by combining multiple (real-valued) filtering functions. Thus, we end this subsection with a few examples of real-valued maps which may be combined to obtain a multifiltering function:
@@
  • Let $K$ be the Delaunay triangulation of a set of points $K_{0}\subset{\mathbb{R}}^{n}$ and consider its filtration in alpha complexes $\{\mathrm{Alpha}_{r}(K_{0})\}_{r\in{\mathbb{R}}}$. Then, for every $\sigma\in K$, note $r_{\sigma}$ the minimal radius parameter $r$ such that $\sigma\in\mathrm{Alpha}_{r}(K_{0})$. The sublevel set $K(r)$ of the radius map $\sigma\mapsto r_{\sigma}$ is exactly $\mathrm{Alpha}_{r}(K_{0})$ for all $r\in{\mathbb{R}}$, meaning that we can see the filtration $\{\mathrm{Alpha}_{r}(K_{0})\}_{r\in{\mathbb{R}}}$ as the sublevel set of the radius map $\sigma\mapsto r_{\sigma}$. A similar reasoning is also valid for point sets filtered in Čech and Vietoris-Rips complexes.
@@ 2.4. Combinatorial vector fields
To ease the computation of multipersistent homology, we can use combinatorial vector fields. Intuitively, we can see a combinatorial vector field on a simplicial complex $K$ as a partition of $K$ into singletons and pairs $\tau<\sigma$ such that $\tau$ is a facet of $\sigma$. More formally, a discrete vector field, or a combinatorial vector field, on a simplicial complex $K$ is an injective partial self-map ${\mathcal{V}}:K\nrightarrow K$ such that
@@ 3. $\mathrm{(3)}$
A combinatorial vector field ${\mathcal{V}}$ is said to be acyclic if all its nontrivial ${\mathcal{V}}$-paths do not loop, i.e. for any sequence of simplices
$\tau_{0}^{(p)},\sigma_{0}^{(p+1)},\tau_{1}^{(p)},\sigma_{1}^{(p+1)},\tau_{2}^{(p)},...,\sigma_{n-1}^{(p+1)},\tau_{n}^{(p)}$
in $K$ such that $\tau_{i}\in\operatorname{dom}{\mathcal{V}}$, ${\mathcal{V}}(\tau_{i})=\sigma_{i}$ and $\sigma_{i}>\tau_{i+1}\neq\tau_{i}$ for each $i=0,...,n-1$ (where $n\geq 1$), we have $\tau_{n}\neq\tau_{0}$.
@@
Consider a multifiltering function $f:K\rightarrow{\mathbb{R}}^{k}$ on a simplicial complex $K$. We say an acyclic combinatorial vector field ${\mathcal{V}}$ is compatible with $f$ or $f$-compatible if $f(\sigma)=f({\mathcal{V}}(\sigma))$ for all $\sigma\in\operatorname{dom}{\mathcal{V}}$.
@@ 3.1. Main definitions
$\mathcal{H}_{g}(\sigma)$ =$=\left\{\gamma^{(p+1)}>\sigma\ |\ g(\gamma)\preceq g(\sigma)\right\};$
@@ 3. $\mathrm{(3)}$
    For all cofacet $\gamma^{(p+1)}>\sigma$, either $g(\gamma)\preceq g(\sigma)$ or $g(\gamma)\succneqq g(\sigma)$.
@@ 4. $\mathrm{(4)}$
${\mathcal{V}}(\sigma)=\begin{cases}\sigma&\text{ if $\sigma$ is critical},\\ \gamma&\text{ if }\mathcal{H}_{g}(\sigma)=\{\gamma\}\text{ for some }\gamma>\sigma,\\ \text{undefined}&\text{ if }\mathcal{T}_{g}(\sigma)\neq\emptyset.\end{cases}$
@@
Notice that for a mdm function $g$, we have $\mathcal{H}_{g}(\sigma)=\{\gamma\}$ iff $\mathcal{T}_{g}(\gamma)=\{\sigma\}$. Hence, the gradient field ${\mathcal{V}}$ of $g$ is such that ${\mathcal{V}}(\sigma)=\gamma$ iff ${\mathcal{V}}(\gamma)$ is undefined and ${\mathcal{V}}^{-1}(\gamma)=\sigma$, and the critical simplices of $g$ are exactly the fixed points ${\mathcal{V}}$. Thus, we often refer to the fixed points of a gradient field ${\mathcal{V}}$ as the critical simplices of ${\mathcal{V}}$ instead. Moreover, we know that a discrete vector field is acyclic if and only if it is the gradient field of some mdm function [14, Proposition 4.9].
Furthermore, it was shown in [4] that if $f:K\rightarrow{\mathbb{R}}^{k}$ is an admissible map which is the max-extension of a component-wise injective vertex map, there exists a $f$-compatible mdm function. One of the main result of this paper, Theorem 4.2, extends this result. Indeed, we will show that for any admissible map $f:K\rightarrow{\mathbb{R}}^{k}$, we can algorithmically define a mdm function $g:K\rightarrow{\mathbb{R}}^{k}$ which, in addition to being $f$-compatible, is obtained from small perturbations of $f$, and is therefore as close as desired to the initial input.
@@ 3.2. Morse inequalities and (relative) perfectness
When the equality holds for each $u\in f(K)$ and each $p\in{\mathbb{Z}}$, we say ${\mathcal{V}}$ is relative-perfect or perfect relatively to $f$. The concept was first introduced in [31] for real-valued filtering functions, while related inequalities have been proved in [34] thereafter.
@@
Consider the triangulated circle $K$ embedded in ${\mathbb{R}}^{2}$ as in Figure 3 and let $f:K\rightarrow{\mathbb{R}}^{2}$ be the max-extension of the map which associates each vertex to its coordinates $(x,y)\in{\mathbb{R}}^{2}$. In this particular case, $f$ happens to be a mdm function, and its gradient field is as in Figure 3. The Betti numbers of the circle are $\beta_{0}(K)=\beta_{1}(K)=1$, so $f$ has too many critical simplices to be a perfect mdm function. Nonetheless, it is relative-perfect. Indeed, if $u\in{\mathbb{R}}^{k}$ is such that $m_{p}(u)=0$ for each $p$, meaning that $L_{u}$ does not contain any critical simplex, then we trivially have the equality $m_{p}(u)=\operatorname{rank}H_{p}\left(K(u),\bigcup_{u^{\prime}\precneqq u}K(u^{\prime})\right)=0$. Otherwise, if $u\in{\mathbb{R}}^{k}$ is such that $m_{p}(u)>0$ for some $p$, we can check that $L_{u}$ contains a single critical simplex $\sigma$ of index $p\in\{0,1\}$, i.e. either a critical vertex or a critical edge. We could then compute that $H_{p}\left(K(u),\bigcup_{u^{\prime}\precneqq u}K(u^{\prime})\right)\cong{\mathbb{Z}}$, so $m_{p}(u)=\operatorname{rank}H_{p}\left(K(u),\bigcup_{u^{\prime}\precneqq u}K(u^{\prime})\right)=1$.